# Products, Abstractions and Inclusions of Causal Spaces

### v1

@@ Abstract
[…]
@@ 1.1 Related Works
Seminal works on causal abstraction with SCMs are Rubenstein et al. [2017] and Beckers and Halpern [2019], where the notions of exact transformations (to be further discussed in Section 5), uniform transformations, abstractions, strong abstractions and constructive abstractions are proposed. Beckers et al. [2020] then relax these to an approximate notion. Massidda et al. [2023] extended the notions to soft interventions, and Zečević et al. [2023] to continually updated abstractions. Causal feature learning is a closely related approach, that also aims to learn higher level features [Chalupka et al., 2015, 2016, 2017]. There are also approaches based on category theory [Rischel and Weichwald, 2021, Otsuka and Saigo, 2022, 2023] and probabilistic logic [Ibeling and Icard, 2023], all grounded in SCMs; see [Zennaro, 2022] for a review.
@@ 2 Preliminaries & Notations
We take a probability space as a starting point, namely, a triple $(\Omega,\mathscr{H},\mathbb{P})$; for a comprehensive introduction, see, for example, [Cinlar, 2011, Durrett, 2019].
Following [Park et al., 2023], we additionally insist that $\mathbb{P}$ is defined over the product measurable space $(\Omega,\mathscr{H})=\otimes_{t\in T}(E_{t},\mathscr{E}_{t})$ with $(E_{t},\mathscr{E}_{t})$ being the same standard measurable space if $T$ is uncountable. Denote by $\mathscr{P}(T)$ the power set of $T$, and for $S\in\mathscr{P}(T)$, we denote by $\mathscr{H}_{S}$ the sub-$\sigma$-algebra of $\mathscr{H}=\otimes_{t\in T}\mathscr{E}_{t}$ generated by measurable rectangles ×$\times_{t\in T}A_{t}$, where $A_{t}\in\mathscr{E}_{t}$ differs from $E_{t}$ only for $t\in S$ and finitely many $t$. In particular, =$\mathscr{H}_{\emptyset}=\{\emptyset,\Omega\}$ is the trivial sub-$\sigma$-algebra of $\Omega=\times_{t\in T}E_{t}$. Also, we denote by $\Omega_{S}$ the subspace ×$\times_{s\in S}E_{s}$ of $\Omega=\times_{t\in T}E_{t}$, and for $T\supseteq S\supseteq U$, we let $\pi_{SU}$ denote the natural projection from $\Omega_{S}$ onto $\Omega_{U}$ and we use the shorthand $\pi_{S}=\pi_{TS}$. We also write $\omega_{S}=\pi_{S}\omega=(\omega_{s})_{s\in S}$.
We write $[d]=\{1,\ldots,d\}$ and $f_{\ast}\mathbb{P}$ denotes the pushforward measure along the map $f$, namely, $f^{*}\mathbb{P}(A)=\mathbb{P}(f^{-1}(A))$.
@@
The causal kernels $K_{S}$ can be defined equivalently as maps from $(\Omega_{S},\mathscr{H}_{S})$ to $(\Omega,\mathscr{H})$, and it will be convenient to use this viewpoint occasionally in our paper (this was also used implicitly in Park et al. [2023]). This observation is explained in Appendix C.
@@
We also recall the definition of causal effects.
@@
Let $(\Omega,\mathscr{H},\mathbb{P},\mathbb{K})=(\times_{t\in T}E_{t},\otimes_{t\in T}\mathscr{E}_{t},\mathbb{P},\mathbb{K})$ be a causal space, $U\subseteq T$ a subset, $A\in\mathscr{H}$ an event and $\mathscr{F}$ a sub-$\sigma$-algebra of $\mathscr{H}$.
We say that $\mathscr{H}_{U}$ is a (local) source of $A$ if $K_{U}(\cdot,A)$ is a version of the conditional probability $\mathbb{P}_{\mathscr{H}_{U}}(A)$.
We say that $\mathscr{H}_{U}$ is a (local) source of $\mathscr{F}$ if $\mathscr{H}_{U}$ is a source of all $A\in\mathscr{F}$. We say that $\mathscr{H}_{U}$ is a global source of the causal space if $\mathscr{H}_{U}$ is a source of all $A\in\mathscr{H}$.
@@ 3 Product Causal Spaces and Causal Independence
Product causal spaces are analogous to connected components in graphical models – see, for example, [Sadeghi and Soo, 2023]. When forming a product of causal spaces that each arise from an SCM, then each component in the product would be a connected component, and the components would not have any causal effect on each other.
@@ 3.1 Causal Independence
The concept of (conditional) independence is arguably one of the most important in probability theory. Now we give an analogous definition of causal independence.
@@
More generally, we say that a finite collection of $\sigma$-algebras $\mathscr{F}_{1},...,\mathscr{F}_{n}$ are a causal independency on $\mathscr{H}_{U}$ if, for all $A_{1}\in\mathscr{F}_{1},...,A_{n}\in\mathscr{F}_{n}$, we have
$K_{U}(\omega,A_{1}\cap...\cap A_{n})=K_{U}(\omega,A_{1})...K_{U}(\omega,A_{n}).$
Moreover, if $I$ is an arbitrary index set, and $\mathscr{F}_{i}$ is a sub-$\sigma$-algebra of $\mathscr{H}$ for each $i\in I$, then the collection $\{\mathscr{F}_{i}:i\in I\}$ is called a causal independency given $\mathscr{H}_{U}$ if its every finite subset is a causal independency on $\mathscr{H}_{U}$.
@@
We use the language of SCMs, because they are convenient framework that fits into the framework of causal spaces.
@@ 4 Transformations of Causal Spaces
Interventional consistency requires that interventions and causal transformations commute, i.e., the result of first intervening and then applying the transformation is the same as intervening on the target after the transformation – see Figure 2.
We emphasise that in Definition 4.1 and 4.2 we do not prescribe conditions for added components indexed by $T^{2}\setminus\rho(T^{1})$.
Further, we remark that as a special case, we can accommodate deterministic maps $f:\Omega_{1}\to\Omega_{2}$ by considering the associated probability kernel $\kappa_{f}(\omega,A)=\boldsymbol{1}_{A}(f(\omega))$. In this case, the admissibility condition reduces to the statement that $\pi_{S}\circ f$ is measurable with respect to $\mathscr{H}^{1}_{\rho^{-1}(S)}$ for all $S\subset\rho(T^{1})$ and distributional consistency becomes, for $A\in\mathscr{H}^{2}$,
@@ 4.1 Examples
Let us provide four prototypical examples of maps between causal spaces that are covered by this definition. Again, we make use of the language of SCMs.
@@ 4.2 Abstractions
We finally remark that one further setting of potential interest would be to consider the inverse of an abstraction, i.e., a setting where a summary variable $X$ is mapped to a more detailed description $(X_{1},X_{2})$. However, to accommodate such transformations we need a slightly different framework than the one presented here. Roughly, we need to consider $\rho:T^{1}\to\mathscr{P}(T^{2})$ with $\rho(t_{1})\cap\rho(t^{\prime}_{1})=\emptyset$ for $t_{1},t^{\prime}_{1}\in T_{1}$, and interventions on all sets $S\subset T^{1}$ can be expressed as interventions on the target $\mathscr{C}^{2}$ (i.e., the more fine-grained representations), while this is reversed in our case so that those two settings are dual to each other.
We do not pursue this here any further, as those transformations are of more limited interest and applicability. Let us emphasise nevertheless that it seems ambitious to handle all cases in one framework. Indeed, combining variables in a summary variable or splitting variables in a more fine-grained description are meaningful operations, but it is less clear to interpret in a causal manner a definition of a transformation $(X_{1},X_{2})\to(Y_{1},Y_{2})$ that allows both at the same time. For example, intervening on $X_{1}$, in general, then does not correspond to a meaningful causal operation on the variables $(Y_{1},Y_{2})$. We also remark that this attempt has not been made in the SCM literature, where the focus is almost exclusively on abstractions.
@@ 5 Comparison with Abstraction in the SCM framework
Finally, we mention that Beckers and Halpern [2019] criticise exact transformations of Rubenstein et al. [2017] on the basis that probabilities and allowed interventions can mask significant differences between SCMs, and then proceed to propose definitions of abstractions that depend only on the structural equations, independently of probabilities. We remark that this criticism is not valid in our framework, in that the interventional consistency of our transformations is imposed independently of probabilities, making it impossible to mask them with the choice of probability measures. That this is possible with SCMs is an artifact of the fact that in SCMs, the observational and interventional measures are coupled through the exogenous distribution, whereas in causal spaces they are completely decoupled.
Moreover, we consider all possible interventions rather than a reduced set of allowed interventions. We also remark that, since probabilities and causal kernels are the primitive objects in our framework, rather than being derived by other primitive objects (namely the structural equations), it does not make sense for the transformation to be defined independently of probabilities, as done by Beckers and Halpern [2019].
@@ 6.1 Existence and Uniqueness of Causal Transformations
Let ${\mathscr{C}}^{2}=(\Omega^{2},\mathscr{H}^{2},\mathbb{P}^{2},{\mathbb{K}}^{2})$ and $\tilde{\mathscr{C}}^{2}=(\Omega^{2},\mathscr{H}^{2},\tilde{\mathbb{P}}^{2},\tilde{\mathbb{K}}^{2})$ be two causal spaces with the same underlying measurable space.
Let $(\kappa,\rho)$ be an admissible pair for the measurable spaces $(\Omega^{1},\mathscr{H}^{1})$ and $(\Omega^{2},\mathscr{H}^{2})$. Assume that the pair $(\kappa,\rho)$ defines causal transformations $\varphi:\mathscr{C}^{1}\to\mathscr{C}^{2}$ and $\tilde{\varphi}:\mathscr{C}^{1}\to\tilde{\mathscr{C}}^{2}$ be a causal transformations.
Then $\mathbb{P}^{2}=\tilde{\mathbb{P}}^{2}$, and for all $A\in\mathscr{H}^{2}_{\rho(T^{1})}$ and any $S\subseteq T^{2}$
$K^{2}_{S}(\omega,A)=\tilde{K}^{2}_{S}(\omega,A)\qquad\text{for }\mathbb{P}^{2}=\tilde{\mathbb{P}}^{2}\text{-a. e. }\omega\in\Omega^{2}.$
The proof is in Appendix B.2.
We cannot expect to derive much stronger results for general causal transformation because interventional consistency does not restrict $K^{2}(\omega,A)$ for $\omega$ not in the support of $\mathbb{P}^{2}$ or $A\notin\mathscr{H}_{\rho(T)}^{2}$. For example, in the setting of Example 4.4, the causal structure on the second factor is arbitrary.
@@
The proof is in Appendix B.2.
To motivate the measurability condition for $K^{1}(\cdot,A)$, we remark that interventional consistency requires $K^{1}(\omega,A)=K^{1}(\omega^{\prime},A)$ for $\omega$ and $\omega^{\prime}$ with $f(\omega)=f(\omega^{\prime})$, and the measurability condition in the result is a slightly stronger condition than this.
@@ 6.2 Sources and causal effects under causal transformations
We now study whether causal effects in target and domain of a causal transformation can be related. Our first results shows that, for perfect abstractions having no causal effect in the domain implies, there is also no causal effect in the target.
@@
The reverse statements are not true, i.e., if there is no causal effect in the target there might be a causal effect in the domain, and if there is an active causal effect in the domain this does not imply that there is a causal effect in the target, which can be seen by considering a target space with only a single point.
We can also study causal effects in the context of embedding transformations, as in Lemma 6.3. Then we see directly that active causal effects are preserved. On the other hand, it is straightforward to construct examples where there is no causal effect in a subsystem, but there is a causal effect in a larger system. This can be achieved by a violation of faithfulness.
@@ Appendix A Background on Probability Theory
Let $T:(\Omega^{1},\mathscr{H}^{1})\to(\Omega^{2},\mathscr{H}^{2})$ be a function. A function $f:\Omega^{1}\to\mathbb{R}$ is $\sigma(T)-\mathscr{B}(\mathbb{R})$ measurable if and only if there is a measurable function $g:(\Omega^{2},\mathscr{H}^{2})\to(\mathbb{R},\mathscr{B}(\mathbb{R}))$ such that $f=g\circ T$.
@@ B.1 Proofs for Section 3
    Here we used that $K_{S^{1}}^{1}(\omega,\Omega_{1})=1=K_{\emptyset}^{1}(\omega,\Omega_{1})$ because $K(\omega,\cdot)$ is a probability measure for a probability kernel.
    So $\mathscr{H}_{T^{1}}$ has no causal effect on $A$. Implication in the other direction follows the same argument.
@@
But by the marginalisation condition on the causal mechanisms $\mathbb{K}$ and $\mathbb{L}$, we have that $L_{S^{\prime}}(\omega,A)=K_{S^{\prime}}(\omega,A)$ for all $\omega\in\Omega^{1}$. This proves interventional consistency. ∎
@@
$\mathbb{P}^{2}(A)=\int\mathbb{P}^{1}(\mathrm{d}\omega)\kappa(\omega,A)=\tilde{P}^{2}(A)$
and thus $\mathbb{P}^{2}=\tilde{\mathbb{P}}^{2}$.
Next, we consider $A\in\mathscr{H}^{2}_{\rho(T^{1})}$ and $S\subset\rho(T^{1})$. Let us define
$B=\{\omega\in\Omega^{2}:K^{2}_{S}(\omega,A)<\tilde{K}_{S}^{2}(\omega,A)\}.$
Since $K^{2}_{S}(\cdot,A)$ and $\tilde{K}^{2}_{S}(\cdot,A)$ are $\mathscr{H}^{2}_{S}$ measurable, we find that $B\in\mathscr{H}^{2}_{S}\subset\mathscr{H}^{2}_{\rho(T^{1})}$. Then the definition of causal spaces (see Definition 2.1) implies that
$K^{2}_{S}(\omega^{\prime},A\cap B)=\boldsymbol{1}_{B}(\omega^{\prime})K^{2}_{S}(\omega^{\prime},A).$
Note that $A\cap B\in\mathscr{H}^{2}_{\rho(T^{1})}$, so we can apply interventional consistency (2) for $\mathscr{C}^{2}$ and $\tilde{\mathscr{C}}^{2}$ and obtain, for any $\omega$,
$\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})K^{2}_{S}(\omega^{\prime},A)$ $=\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})K^{2}_{S}(\omega^{\prime},A\cap B)$
$=\int K^{1}_{\rho^{-1}(S)}(\omega,\mathrm{d}\omega^{\prime})\kappa(\omega^{\prime},A)$
$=\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})\tilde{K}^{2}_{S}(\omega^{\prime},A\cap B)$
$=\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})\tilde{K}^{2}_{S}(\omega^{\prime},A).$
@@
00 0 $=\int\mathbb{P}^{1}(\mathrm{d}\omega)\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})(\tilde{K}^{2}_{S}(\omega^{\prime},A)-K^{2}_{S}(\omega^{\prime},A))$
$=\int\mathbb{P}^{2}(\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})(\tilde{K}^{2}_{S}(\omega^{\prime},A)-K^{2}_{S}(\omega^{\prime},A))$
$=\int_{B}\mathbb{P}^{2}(\mathrm{d}\omega^{\prime})\,(\tilde{K}^{2}_{S}(\omega^{\prime},A)-K^{2}_{S}(\omega^{\prime},A)).$
On $B$, the integrand is strictly positive by definition. Thus we conclude that $\mathbb{P}^{2}(B)=0$ and thus $\tilde{K}^{2}_{S}(\omega^{\prime},A)\leq{K}^{2}_{S}(\omega^{\prime},A)$ holds almost surely.
The same reasoning implies the reverse bound, and we conclude that, $\mathbb{P}^{2}$-almost surely, the relation
$\tilde{K}^{2}_{S}(\omega^{\prime},A)={K}^{2}_{S}(\omega^{\prime},A)$

### v2

@@ Abstract
Causal spaces have recently been introduced as a measure-theoretic framework to encode the notion of causality. While it has some advantages over established frameworks, such as structural causal models, the theory is so far only developed for single causal spaces. In many mathematical theories, not least the theory of probability spaces of which causal spaces are a direct extension, combinations of objects and maps between objects form a central part. In this paper, taking inspiration from such objects in probability theory, we propose the definitions of products of causal spaces, as well as (stochastic) transformations between causal spaces. In the context of causality, these quantities can be given direct semantic interpretations as causally independent components, abstractions and extensions.
@@ 1.1 Related Works
Seminal works on causal abstraction with SCMs are Rubenstein et al. [2017] and Beckers and Halpern [2019], where the notions of exact transformations (to be further discussed in Section 5), uniform transformations, abstractions, strong abstractions and constructive abstractions are proposed. Beckers et al. [2020] then relax these to an approximate notion. Massidda et al. [2023] extended the notions to soft interventions, and Zečević et al. [2023] to continually updated abstractions. Causal feature learning is a closely related approach, that also aims to learn higher level features [Chalupka et al., 2015, 2016, 2017] There are also approaches based on category theory [Rischel and Weichwald, 2021, Otsuka and Saigo, 2022, 2023] and probabilistic logic [Ibeling and Icard, 2023], all grounded in SCMs; see [Zennaro, 2022] for a review.
@@ 2 Preliminaries & Notations
We take a probability space as a starting point, namely, a triple $(\Omega,\mathscr{H},\mathbb{P})$; for a comprehensive introduction, see, for example, [Cinlar, 2011, Durrett, 2019]. Following [Park et al., 2023], we additionally insist that $\mathbb{P}$ is defined over the product measurable space $(\Omega,\mathscr{H})=\otimes_{t\in T}(E_{t},\mathscr{E}_{t})$ with $(E_{t},\mathscr{E}_{t})$ being the same standard measurable space if $T$ is uncountable. Denote by $\mathscr{P}(T)$ the power set of $T$, and for $S\in\mathscr{P}(T)$, we denote by $\mathscr{H}_{S}$ the sub-$\sigma$-algebra of $\mathscr{H}=\otimes_{t\in T}\mathscr{E}_{t}$ generated by measurable rectangles ×$\times_{t\in T}A_{t}$, where $A_{t}\in\mathscr{E}_{t}$ differs from $E_{t}$ only for $t\in S$ and finitely many $t$. In particular, =$\mathscr{H}_{\emptyset}=\{\emptyset,\Omega\}$ is the trivial sub-$\sigma$-algebra of $\Omega=\times_{t\in T}E_{t}$. Also, we denote by $\Omega_{S}$ the subspace ×$\times_{s\in S}E_{s}$ of $\Omega=\times_{t\in T}E_{t}$, and for $T\supseteq S\supseteq U$, we let $\pi_{SU}$ denote the natural projection from $\Omega_{S}$ onto $\Omega_{U}$ and we use the shorthand $\pi_{S}=\pi_{TS}$. We also write $\omega_{S}=\pi_{S}\omega=(\omega_{s})_{s\in S}$. We write $[d]=\{1,\ldots,d\}$ and $f_{\ast}\mathbb{P}$ denotes the pushforward measure along the map $f$, namely, $f^{*}\mathbb{P}(A)=\mathbb{P}(f^{-1}(A))$.
@@
The causal kernels $K_{S}$ can be defined equivalently as maps from $(\Omega_{S},\mathscr{H}_{S})$ to $(\Omega,\mathscr{H})$ and it will be convenient to use this viewpoint occasionally in the following (this was also used implicitly in Park et al. [2023]). This observation is explained in Appendix C.
@@
We also recall the definition of causal effect.
@@
Let $(\Omega,\mathscr{H},\mathbb{P},\mathbb{K})=(\times_{t\in T}E_{t},\otimes_{t\in T}\mathscr{E}_{t},\mathbb{P},\mathbb{K})$ be a causal space, $U\subseteq T$ a subset, $A\in\mathscr{H}$ an event and $\mathscr{F}$ a sub-$\sigma$-algebra of $\mathscr{H}$. We say that $\mathscr{H}_{U}$ is a (local) source of $A$ if $K_{U}(\cdot,A)$ is a version of the conditional probability $\mathbb{P}_{\mathscr{H}_{U}}(A)$. We say that $\mathscr{H}_{U}$ is a (local) source of $\mathscr{F}$ if $\mathscr{H}_{U}$ is a source of all $A\in\mathscr{F}$. We say that $\mathscr{H}_{U}$ is a global source of the causal space if $\mathscr{H}_{U}$ is a source of all $A\in\mathscr{H}$.
@@ 3 Product Causal Spaces and Causal Independence
Product causal spaces are analogous to connected components in graphical models – see, for example, [Sadeghi and Soo, 2023].
@@ 3.1 Causal Independence
Now we give a definition of causal independence that is analogous to conditional independence.
@@
Again, causal independence is only defined for two $\sigma$-algebras for the sake of notational convenience; it can easily be extended to arbitrary collections of $\sigma$-algebra.
@@ 4 Transformations of Causal Spaces
Interventional consistency requires that interventions and causal transformations commute, i.e., the result of first intervening and then applying the transformation is the same as intervening on the target after the transformation – see Figure 2. We emphasise that in Definition 4.1 and 4.2 we do not prescribe conditions for added components indexed by $T^{2}\setminus\rho(T^{1})$. Further, we remark that as a special case, we can accommodate deterministic maps $f:\Omega_{1}\to\Omega_{2}$ by considering the associated probability kernel $\kappa_{f}(\omega,A)=\boldsymbol{1}_{A}(f(\omega))$. In this case, the admissibility condition reduces to the statement that $\pi_{S}\circ f$ is measurable with respect to $\mathscr{H}^{1}_{\rho^{-1}(S)}$ for all $S\subset\rho(T^{1})$ and distributional consistency becomes, for $A\in\mathscr{H}^{2}$,
@@ 4.1 Examples
Let us provide four prototypical examples of maps between causal spaces that are covered by this definition. Here we resort to the language of SCMs because they are a convenient framework that fits into causal spaces.
@@ 4.2 Abstractions
We finally remark that one further setting of potential interest would be to consider the inverse of an abstraction, i.e., a setting where a summary variable $X$ is mapped to a more detailed description $(X_{1},X_{2})$. However, to accommodate such transformations we need a slightly different framework than the one presented here. Roughly, we need to consider $\rho:T^{1}\to\mathscr{P}(T^{2})$ with $\rho(t_{1})\cap\rho(t^{\prime}_{1})=\emptyset$ for $t_{1},t^{\prime}_{1}\in T_{1}$, and interventions on all sets $S\subset T^{1}$ can be expressed as interventions on the target $\mathscr{C}^{2}$ (i.e., the more fine-grained representations), while this is reversed in our case so that those two settings are dual to each other. We do not pursue this here any further, as those transformations are of more limited interest and applicability. Let us emphasise nevertheless that it seems ambitious to handle all cases in one framework. Indeed, combining variables in a summary variable or splitting variables in a more fine-grained description are meaningful operations, but it is less clear to interpret in a causal manner a definition of a transformation $(X_{1},X_{2})\to(Y_{1},Y_{2})$ that allows both at the same time. For example, intervening on $X_{1}$, in general, then does not correspond to a meaningful causal operation on the variables $(Y_{1},Y_{2})$. We also remark that this attempt has not been made in the SCM literature, where the focus is almost exclusively on abstractions.
@@ 5 Comparison with Abstraction in the SCM framework
Finally, we mention that Beckers and Halpern [2019] criticise exact transformations of Rubenstein et al. [2017] on the basis that probabilities and allowed interventions can mask significant differences between SCMs, and then proceed to propose definitions of abstractions that depend only on the structural equations, independently of probabilities. We remark that this criticism is not valid in our framework, in that the interventional consistency of our transformations is imposed independently of probabilities, making it impossible to mask them with the choice of probability measures. That this is possible with SCMs is an artifact of the fact that in SCMs, the observational and interventional measures are coupled through the exogenous distribution, whereas in causal spaces they are completely decoupled. Moreover, we consider all possible interventions rather than a reduced set of allowed interventions. We also remark that, since probabilities and causal kernels are the primitive objects in our framework, rather than being derived by other primitive objects (namely the structural equations), it does not make sense for the transformation to be defined independently of probabilities, as done by Beckers and Halpern [2019].
@@ 6.1 Existence and Uniqueness of Causal Transformations
Let ${\mathscr{C}}^{2}=(\Omega^{2},\mathscr{H}^{2},\mathbb{P}^{2},{\mathbb{K}}^{2})$ and $\tilde{\mathscr{C}}^{2}=(\Omega^{2},\mathscr{H}^{2},\tilde{\mathbb{P}}^{2},\tilde{\mathbb{K}}^{2})$ be two causal spaces with the same underlying measurable space. Let $(\kappa,\rho)$ be an admissible pair for the measurable spaces $(\Omega^{1},\mathscr{H}^{1})$ and $(\Omega^{2},\mathscr{H}^{2})$. Assume that the pair $(\kappa,\rho)$ defines causal transformations $\varphi:\mathscr{C}^{1}\to\mathscr{C}^{2}$ and $\tilde{\varphi}:\mathscr{C}^{1}\to\tilde{\mathscr{C}}^{2}$ be a causal transformations. Then $\mathbb{P}^{2}=\tilde{\mathbb{P}}^{2}$, and for all $A\in\mathscr{H}^{2}_{\rho(T^{1})}$ and any $S\subset T^{2}$
$K^{2}_{S}(\omega,A)=\tilde{K}^{2}_{S}(\omega,A)\quad\text{for $\mathbb{P}^{2}=\tilde{\mathbb{P}}^{2}$ a.\ e.\ $\omega\in\Omega^{2}$}.$
The proof is in Appendix B.2. We cannot expect to derive much stronger results for general causal transformation because interventional consistency does not restrict $K^{2}(\omega,A)$ for $\omega$ not in the support of $\mathbb{P}^{2}$ or $A\notin\mathscr{H}_{\rho(T)}^{2}$. E.g., in the setting of Example 4.4 the causal structure on the second factor is arbitrary.
@@
The proof is in Appendix B.2. To motivate the measurability condition for $K^{1}(\cdot,A)$ we remark that interventional consistency requires $K^{1}(\omega,A)=K^{1}(\omega^{\prime},A)$ for $\omega$, $\omega^{\prime}$ with $f(\omega)=f(\omega^{\prime})$ and the measurability condition in the result is a slightly stronger condition than this.
@@ 6.2 Sources and causal effects under causal transformations
We now study whether causal effects in target and domain of a causal transformation can be related. Our first results shows that for perfect abstractions having no causal effect in the domain implies that there is also no causal effect in the target.
@@
The reverse statements are not true, i.e., if there is no causal effect in the target there might be a causal effect in the domain and if there is an active causal effect in the domain this does not imply that ther is a causal effect in the target, which can be seen by considering a target space with only a single point. We can also study causal effects in the context of embedding transformations, as in Lemma 6.3. Then we see directly that active causal effects are preserved. On the other hand, it is straightforward to construct examples where there is no causal effect in a subsystem, but there is a causal effect in a larger system. This can be achieved by a violation of faithfulness.
@@ 7 Conclusion
Acknowledgements.
This work was supported by the Tübingen AI Center.
@@ Appendix A Background on Probability Theory
Let $T:\Omega^{1}\to(\Omega^{2},\mathscr{H}^{2})$ be a function. A function $f:\Omega^{1}\to\mathbb{R}$ is $\sigma(T)$$\mathscr{B}(\mathbb{R})$ measurable if and only if there is a measurable function $g:(\Omega^{2},\mathscr{H}^{2})\to(\mathbb{R},\mathscr{B}(\mathbb{R}))$ such that $f=g\circ T$.
@@ B.1 Proofs for Section 3
    Here we used that $K_{S^{1}}^{1}(\omega,\Omega_{1})=1=K_{\emptyset}^{1}(\omega,\Omega_{1})$ because $K(\omega,\cdot)$ is a probability measure for a probability kernel. So $\mathscr{H}_{T^{1}}$ has no causal effect on $A$. Implication in the other direction follows the same argument.
@@
But by the marginalisation condition on the causal mechanisms $\mathbb{K}$ and $\mathbb{L}$, we have that $L_{S^{\prime}}(\omega,A)=K_{S^{\prime}}(\omega,A)$ for all $\omega\in\Omega^{1}$. This proves interventional consistency.
@@
$\mathbb{P}^{2}(A)=\int\mathbb{P}^{1}(\mathrm{d}\omega)\kappa(\omega,A)=\tilde{P}^{2}(A)$ (4)
and thus $\mathbb{P}^{2}=\tilde{\mathbb{P}}^{2}$. Next, we consider $A\in\mathscr{H}^{2}_{\rho(T^{1})}$ and $S\subset\rho(T^{1})$. Let us define
$B=\{\omega\in\Omega^{2}:K^{2}_{S}(\omega,A)<\tilde{K}_{S}^{2}(\omega,A)\}.$ (5)
Since $K^{2}_{S}(\cdot,A)$ and $\tilde{K}^{2}_{S}(\cdot,A)$ are $\mathscr{H}^{2}_{S}$ measurable we find $B\in\mathscr{H}^{2}_{S}\subset\mathscr{H}^{2}_{\rho(T^{1})}$. Then the definition of causal spaces (see Definition 2.1) implies that
$K^{2}_{S}(\omega^{\prime},A\cap B)=\boldsymbol{1}_{B}(\omega^{\prime})K^{2}_{S}(\omega^{\prime},A).$ (6)
Note that $A\cap B\in\mathscr{H}^{2}_{\rho(T^{1})}$ so we can apply interventional consistency (2) for $\mathscr{C}^{2}$ and $\tilde{\mathscr{C}}^{2}$ and obtain for any $\omega$
$\begin{split}\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})K^{2}_{S}(\omega^{\prime},A)&=\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})K^{2}_{S}(\omega^{\prime},A\cap B)=\int K^{1}_{\rho^{-1}(S)}(\omega,\mathrm{d}\omega^{\prime})\kappa(\omega^{\prime},A)\\ &=\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})\tilde{K}^{2}_{S}(\omega^{\prime},A\cap B)=\int\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})\tilde{K}^{2}_{S}(\omega^{\prime},A)\end{split}$ (7)
@@
$\begin{split}0&=\int\mathbb{P}^{1}(\mathrm{d}\omega)\kappa(\omega,\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})(\tilde{K}^{2}_{S}(\omega^{\prime},A)-K^{2}_{S}(\omega^{\prime},A))\\ &=\int\mathbb{P}^{2}(\mathrm{d}\omega^{\prime})\boldsymbol{1}_{B}(\omega^{\prime})(\tilde{K}^{2}_{S}(\omega^{\prime},A)-K^{2}_{S}(\omega^{\prime},A))\\ &=\int_{B}\mathbb{P}^{2}(\mathrm{d}\omega^{\prime})\,(\tilde{K}^{2}_{S}(\omega^{\prime},A)-K^{2}_{S}(\omega^{\prime},A)).\end{split}$ (8)
On $B$ the last term is strictly positive by definition. Thus we conclude that $\mathbb{P}^{2}(B)=0$ and thus $\tilde{K}^{2}_{S}(\omega^{\prime},A)\leq{K}^{2}_{S}(\omega^{\prime},A)$ holds almost surely. The same reasoning implies the reverse bound and we conclude that $\mathbb{P}^{2}$ almost surely the relation
$\tilde{K}^{2}_{S}(\omega^{\prime},A)={K}^{2}_{S}(\omega^{\prime},A)$ (9)